. Your firm recently sold an asset and financed the sale of the asset. The firm took back a $170,000 loan on the equipment sale with monthly payments and 8% interest. The loan was a 5-year loan. Your firm now needs cash and desires to sell the loan. If investors require 12% rate of return to invest in this type of loan and 6 payments have been made on the loan, how much will your firm loose from the sale of this financial asset?
Answer #1

First step:

Existing monthly payments:

Monthly payment = [P × R × (1+R)^N ] / [(1+R)^N -1]
Using the formula:
Loan amount P $                                                          170,000
Rate of interest per period:
Annual rate of interest 8.000%
Frequency of payment = Once in 1 month period
Numer of payments in a year = 12/1 = 12
Rate of interest per period R 0.08 /12 = 0.6667%
Total number of payments:
Frequency of payment = Once in 1 month period
Number of years of loan repayment =                                                                          5
Total number of payments N 5 × 12 = 60
Period payment using the formula = [ 170000 × 0.00667 × (1+0.00667)^60] / [(1+0.00667 ^60 -1]
Monthly payment = $                                                         3,446.99

Second step:

Balance in loan principal after six payments:

Loan balance = PV * (1+r)^n - P[(1+r)^n-1]/r
Loan amount PV = 170,000.00
Rate of interest r= 0.6667%
nth payment n= 6
Payment P= 3,446.99
Loan balance = 170000*(1+0.00667)^6 - 3446.99*[(1+0.00667)^6-1]/0.00667
Loan balance =                                                                         155,884.63

Existing loan balance is $155,884.63

Third step:

Sale value of loan with existing monthly payments for 54 months is:

a Present value of annuity= P* [ [1- (1+r)-n ]/r ]
P= Periodic payment $                  3,446.99
r= Rate of interest per period
Annual interest 12.00%
Number of interest payments per year 12
Interest rate per period 0.12/12=
Interest rate per period 1.000%
n= number of periods:
Number of years 4.5
Periods per year 12
number of payments 54
Present value of annuity= 3446.99* [ (1- (1+0.01)^-54)/0.01 ]
Present value of annuity= 143,286.77

Loss on sale of loan = $155,884.63 - $143,286.77 = $12,597.86

Please rate.